solve for the variable in each equation below. show all work. check answer. if a solution is undefined, state that in answer.
3/a+1=5/a+7
please help if you can. thanks

Answer by stanbon(75887)

solve for the variable in each equation below. show all work. check answer. if a solution is undefined, state that in answer.
3/a+1=5/a+7
----
Cross-multiply:
3(a+7) = 5(a+1)
3a+21 = 5a+5
2a = 16
a = 8
===================
Cheers,
Stan H.
===============